OpenStudy (anonymous):

A carpenter is building a bookcase the perimeter of the top is 108 in. What length must the stock be if the width is 8 in.? The shelf is_____inches long.

OpenStudy (anonymous):

\[Perimeter = 2 \times width + 2 \times length\] \[180in = 2 \times 8in + 2 \times length\] Does this seem correct? Do you know how to isolate a variable?

OpenStudy (anonymous):

L=46